Mary invested 10000 part at 4% and part at 5% for 1 year. How much was invested in each acct if interest earned in the 5% acct was 284 greater than 4% acct. the
Answer by richwmiller(17219)

We know the total amount of money invested. $10000
x+y=10000,
We know that the difference in interest earned by the two accounts is $284
0.05*x-0.04*y=284
x=10000-y
We substitute for x
0.05*(10000-y)-0.04*y=284
We multiply out
500-0.05y-0.04*y=284
We combine like terms.
216=0.09*y
Isolate y
y=216/0.09
y=2400 at 4%
Calculate x
x=10000-2400
x=7600 at 5%
Interest earned at 5% is 380
Interest earned at 4% is 96
We check
0.05*7600-0.04*2400=284
380-96=284
284=284
Since this statement is TRUE and neither x nor y is negative then all is well
codeintdy